OpenStudy (anonymous):

Solve the system of equations by the addition method. 5y+y=23 2x-y=5 PLEASE HELP

OpenStudy (jdoe0001):

hint: " by the addition method. "

OpenStudy (jdoe0001):

\(\large \begin{array}{cllll} 5y+y=23\\ 2x-y=5 \\\hline\\ \square ?+\square ?=\square ? \end{array}\)